OpenStudy (anonymous):

Simplify by factoring ᶟ√250 need help plz

OpenStudy (anonymous):

3952.8

OpenStudy (anonymous):

\[\sqrt[3]{250}=\sqrt[3]{2*5*5*5}=5\sqrt[3]{2} \]